# KACQ-DCNN: Uncertainty-Aware Interpretable Kolmogorov–Arnold Classical–Quantum Dual-Channel Neural Network for Heart Disease Detection

### v3

@@ Methods
Statistical validation
@@
Interpretability analysis
@@
LIME: local interpretability
@@
Figure LABEL:fig:lime_heart_failure shows the importance of and , which are prominent in determining the classification outcome, with contrasting effects between HF and normal classifications. Figure LABEL:fig:lime_normal_heart presents the LIME explanation for predicting a healthy heart condition. Features like between 0.50 and 1.00, absence of (), and values of chest pain type within the range of positively contribute (green bars) to the normal classification. Conversely, certain factors such as and deviations negatively influence the likelihood of a normal classification (red bars). Figure LABEL:fig:lime_heart_failure shows the top contributing features towards the classification of HF include a lower ST_Slope (), absence of (), high (), and (). These features have positive contributions (green bars) to HF prediction, with ST_Slope having the largest impact. Some features, like (), have negative contributions (red bars), indicating their absence reduces the likelihood of HF. This comparison between the two LIME visualizations highlights how the same features can have contrasting effects depending on their values and the predicted class.
SHAP: global interpretability
@@
Applying SHAP to our model provided a global perspective on the importance of features and their contributions to predictions. SHAP decision plot (Figure LABEL:fig:shap_decision_plot) further demonstrated how features such as , , , and significantly influence the prediction, as indicated by the sharp inflection points. These features push the model output towards a higher prediction score. Other features such as and have less impact on the model’s final decision, as shown by more minor deviations in the path. SHAP waterfall plot in Figure LABEL:fig:shap_waterfall_plot visualizes the contribution of various features to a prediction made by the model. The base value , adjusted by each feature’s contribution, represents the average model output, whereas the final model prediction is shown at the top. The most influential features include (+0.12), (+0.1), and (+0.06), which all significantly increase the predicted probability. Conversely, features such as (-0.02) and (-0.01) have negative impacts, slightly reducing the model’s prediction. SHAP force plot in Figure LABEL:fig:shap_force_plot visualizes how individual features influence a specific prediction by showing the combined positive and negative contributions of each feature. For this particular prediction, the most impactful positive features are and , which push the prediction upwards towards 0.84. Conversely, features such as , , and have a slight negative effect, reducing the prediction. The plot effectively highlights which factors led to the final prediction and in which direction they influenced the outcome. Figure LABEL:fig:shap_force_plot2 summarizes multiple predictions, revealing the cumulative effect of features across different samples. For this group of predictions, and emerge as highly influential features. Other features like and also play significant roles, but appears repeatedly as a strong driver in either direction. The red areas show that for many instances, and tend to increase the predicted probability of heart disease, while blue areas for push it down for others. This global view emphasizes the consistency of these critical features in shaping the model’s predictions, reflecting their significant role in cardiac risk assessment.
Conformal prediction analysis
@@
The calibration curve in Figure LABEL:fig:uq1 illustrates the relationship between the predicted probabilities of KACQ-DCNN and the actual fraction of positive outcomes. An ideal calibration curve would follow the diagonal line, indicating that the predicted probabilities align perfectly with the true probabilities. However, as observed, the model exhibits miscalibration in several regions. At lower-mid predicted probabilities (30%–60%), the model tends to underestimate the occurrence of positive outcomes, with the curve falling below the ideal line. Conversely, in the upper-mid range of predicted probabilities (60%–80%), the model overestimates the likelihood of positive outcomes, as indicated by the curve exceeding the diagonal. This overestimation can result in a greater number of FPs. At lower (0%–30%) and higher probabilities (80%–100%), the model’s predictions are closer to the actual outcomes, although slight overprediction remains. These deviations indicate that the predicted probabilities do not fully represent the true likelihoods, especially in the lower-mid and upper-mid ranges. To address this issue, recalibration methods such as Platt Scaling or Isotonic Regression could be employed to improve the alignment between predicted probabilities and actual outcomes.
@@
Figure LABEL:fig:uq3 shows that the Standard conformal predictor (left plot) deviates from the diagonal line, mainly in the middle range of p-values. This suggests that it might not be as confident in its predictions for some ranges of positive class probabilities. Similarly, the Mondrian conformal predictor (right plot) also demonstrates deviations from the ideal line. However, the nature and magnitude of these deviations differ because of the additional conditioning on feature bins (Mondrian approach). The Standard and Mondrian conformal predictors initialized via “crepes” were tested on several important metrics, including error rate, average prediction set size, single-class predictions, and number of empty sets. In addition, we recorded the computational efficiency with respect to fitting and evaluation time. These results provide insights into the calibration, precision, and computational feasibility of the model.
@@
We developed a custom wrapper to integrate a TensorFlow-based sequential model with the MAPIE framework for conformal predictions. Early stopping was applied on the basis of validation loss to prevent overfitting. We generated class probability estimates by leveraging the model’s outputs, which were then used to form class predictions based on a threshold of 0.5. This TensorFlow-wrapped model was then integrated into MAPIE’s conformal prediction framework to estimate prediction sets with a confidence level . Specifically, we use sigmoid outputs from the wrapper model as conformity scores, and MAPIE computes the conformal quantile , which helps define the prediction sets for new data points. These prediction sets were designed to include all classes whose sigmoid outputs exceeded , ensuring that the true label was included in the set with a specified probability. Finally, we evaluated the model via different methods (‘score’ and ‘lac’). We tested multiple confidence levels () to generate prediction intervals, ensuring the inclusion of true labels with high reliability across test datasets. Figure LABEL:fig:uq2 shows that the estimated quantile values decreased as increased. A higher results in a lower quantile threshold, including more data points in the prediction set. In contrast, a lower (e.g., ) leads to a higher quantile threshold, allowing fewer but more confident predictions. Consequently, when is exceptionally high, it may lead to a low quantile value, resulting in large prediction regions. Conversely, if is too small, the quantile increases, potentially leaving no null areas of the prediction set.
@@
As shown in Figure LABEL:fig:uq4, when class coverage is low, there may be times when the model is unsure about its predictions, especially close to decision boundaries. This can cause prediction sets to be empty. As class coverage increases, these empty regions disappear, but areas of ambiguity emerge where both classes may be included in the prediction sets. This means that the selection of class coverage directly influences the prediction decisions and vice versa. A notable situation arises when predictions are based on a threshold of 0.5, as in our binary classification. In this case, each prediction set contained only a single class, eliminating ambiguous or uncertain regions. Figure LABEL:fig:uq6 shows that both methods provide coverage close to the target coverage, regardless of the value. The average size of prediction sets increased, and the number of empty prediction sets decreased sharply as increased. The histograms in Figure LABEL:fig:uq5 help visualize how often each method produces single or multiple class predictions. The bars for size = 0 are insignificant, implying that the methods do not struggle with uncertainty. Similarly, a dominance of size = 1 indicates strong model confidence.
emissions in our research
@@
The experiments were carried out on the Google Cloud Platform, specifically in the “asia-east1” region, which has a carbon efficiency rating of 0.56 kg of equivalent per kilowatt-hour (kWh). The research conducted hours of computation using T4 hardware, which has a thermal design power (TDP) of 70W. The estimated total emissions were 15.68 kg of CO2 equivalent, all of which were fully offset by the cloud service provider. To clarify the calculations, the power consumption was determined by multiplying the power rating (70W) by the computation time (400 hours), resulting in a total energy consumption of 28 kWh. This figure was then multiplied by the region’s carbon efficiency of 0.56 kg eq/kWh to yield the total emissions of 15.68 kg eq. If the same model was executed in the Google Cloud Platform’s “europe-west6” region, the carbon emissions would have been lower, at 0.45 kg of CO2 equivalent per kWh.
@@
Conclusions
In this study, we propose the KACQ-DCNN model, a novel classical-quantum hybrid architecture that replaces MLP with KAN layers, demonstrating promising results in heart disease diagnosis. Our comprehensive benchmarking revealed that the KAN-enhanced KACQ-DCNN 1-Layered model outperformed 34 benchmark models, including 16 CML models, 12 QNN models, and 6 hybrid models. Specifically, our KACQ-DCNN 1-Layered model achieved an accuracy of , a maP of , a maR of , and a maF1 score of . Additional metrics further support the model’s effectiveness, with a ROC-AUC score of , MCC of , and Cohen’s kappa of . The robustness of these results was confirmed via stratified 10-fold cross-validated two-tailed paired t-tests. Statistical significance was established via the Bonferroni correction, with an value of . All -values were substantially lower than this threshold, allowing us to reject the for all models and confirming the superiority of KACQ-DCNN. Our ablation study highlights the synergistic effects of the hybrid approach. The ROC-AUC column shows a significant variance in the ablation test, but the KAN-version of KACQ-DCNN 1-Layered outperforms all the other metrics when every component works in harmony. Although the dual-channel design choice may diminish the architecture’s effectiveness in real-time scenarios, the versatility of the KACQ-DCNN and its capacity to handle complex datasets with high precision makes it an optimal choice for our objectives.
@@ Conclusions
Model-agnostic XAI, such as LIME and SHAP, helps us precisely find the key features that affect predictions. They also improved the KACQ-DCNN model’s local and global interpretability and trustworthiness by validating feature contributions and ensuring that decisions were made consistently across different instances. Uncertainty quantification analysis provided vital insights into the predictive performance and reliability of the KACQ-DCNN model. The calibration curve indicated notable miscalibration in specific probability ranges, where the model underestimated and overestimated positive outcomes. This misalignment suggests that recalibration methods, such as Platt scaling or isotonic regression, could improve the model’s probability estimates. The Mondrian approach slightly enhances reliability with a lower error rate and smaller prediction set sizes. This led to more accurate predictions when the Standard and Mondrian conformal predictors were used for conformal prediction analysis. Additionally, the Mondrian method demonstrated higher confidence by producing single-class predictions in more cases. Further analysis revealed that as the confidence level increased (lower values), the prediction sets became smaller, offering more confident predictions, whereas higher values led to broader prediction regions.
While our findings show promise for enhancing heart disease prediction, they are not without limitations. One potential limitation is the dependency on the availability of quantum hardware, which is still in the early stages of development and may limit its practical implementation and scalability. Qubits can lose entanglement from heat and light, causing quantum decoherence and data loss. Incorrect qubit rotations can lead to inaccurate results, potentially causing misdiagnosis. Current CQML models are limited by these quantum errors, which affect their reliability in sensitive applications. Additionally, the complexity of model training may pose challenges, as optimizing a hybrid architecture requires careful tuning of classical and quantum components to achieve optimal performance. The performance of KACQ-DCNN may be sensitive to data quality and representativeness, particularly given the inherent challenges associated with short and tabular datasets. Future research could explore integrating more diverse datasets to improve generalizability, investigate advanced techniques for uncertainty quantification, and further refine the model architecture to enhance interpretability. Additionally, as quantum computing technology matures, future work could leverage more powerful quantum systems to unlock even more significant computational advantages, and extending KACQ-DCNN to diagnose related conditions such as diabetes and sleep apnea could broaden its utility in medical diagnostics.
@@
The datasets used in this study are publicly available on Kaggle at the following link: https://www.kaggle.com/datasets/fedesoriano/heart-failure-prediction.

### v4

@@ 4.1.3 Ablation study analysis
The key observations from the ablation study (Table 7) are:
1.
The KAN-version of KACQ-DCNN (92.03% accuracy) substantially outperforms the MLP version (90.00% accuracy) with a similar parameter count, underscoring KAN’s superior function approximation capabilities.
2.
Replacing BiLSTMs with LSTMs or removing BiLSTM layers led to a noticeable drop in accuracy (to 89%), indicating the importance of bidirectional sequential processing. However, the model maintained some robustness.
3.
Removing DenseKAN layers from the classical channel reduced accuracy to 88%, highlighting their crucial role.
4.
Removing quantum layers reduced accuracy back to 89% and significantly impacted MCC and Kappa, confirming the positive contribution of the quantum components. Conversely, removing DenseKAN layers from the quantum channel improved accuracy to 91%, suggesting a simplification in that part of the model was beneficial.
5.
Using angle embedding instead of amplitude embedding for quantum circuits decreased accuracy to 88%, suggesting AE is more suitable for this architecture.
6.
Removing dropout layers maintained 90% accuracy and reduced training time, but the top-performing KACQ-DCNN includes them, suggesting they help achieve the peak 92.03% by mitigating overfitting during more extensive optimization.
4.1.4 Statistical validation
@@
Key observations from the statistical validation in Table 8 include:
1.
All calculated p-values for the paired t-tests between KACQ-DCNN 1-Layered and the nine other top-performing models are substantially lower than the Bonferroni-corrected significance level of .
2.
This allows us to confidently reject for all comparisons, indicating a statistically significant difference in accuracy.
3.
Cohen’s d values are consistently large (ranging from 4.310 to 9.558), demonstrating that the observed accuracy improvements of KACQ-DCNN 1-Layered are not only statistically significant but also practically meaningful in magnitude.
4.
These results provide robust statistical evidence that the proposed KACQ-DCNN 1-Layered model significantly improves prediction accuracy over the other benchmarked models for heart disease detection on this dataset.
4.2 Interpretability analysis
@@
4.2.1 LIME: local interpretability
@@
Figure 10(b) shows the importance of and , which are prominent in determining the classification outcome, with contrasting effects between HF and normal classifications. Figure 10(a) presents the LIME explanation for predicting a healthy heart condition. Features like between 0.50 and 1.00, absence of (), and values of chest pain type within the range of positively contribute (green bars) to the normal classification. Conversely, certain factors such as and deviations negatively influence the likelihood of a normal classification (red bars). Figure 10(b) shows that the top contributing features towards the classification of HF include a lower ST_Slope (), absence of (), high (), and (). These features have positive contributions (green bars) to HF prediction, with having the largest impact. Some features, like (), have negative contributions (red bars), indicating their absence reduces the likelihood of HF. This comparison between the two LIME visualizations highlights how the same features can have contrasting effects depending on their values and the predicted class.
4.2.2 SHAP: global interpretability
@@
Applying SHAP to our model provided a global perspective on the importance of features and their contributions to predictions. SHAP decision plot (Figure 11(a)) further demonstrated how features such as , , , and significantly influence the prediction, as indicated by the sharp inflection points. These features push the model output towards a higher prediction score. Other features, such as and have less impact on the model’s final decision, as shown by more minor deviations in the path. SHAP waterfall plot in Figure 11(b) visualizes the contribution of various features to a prediction made by the model. The base value , adjusted by each feature’s contribution, represents the average model output, whereas the final model prediction is shown at the top. The most influential features include (+0.12), (+0.1), and (+0.06), which all significantly increase the predicted probability. Conversely, features such as (-0.02) and (-0.01) have negative impacts, slightly reducing the model’s prediction. SHAP force plot in Figure 12(a) visualizes how individual features influence a specific prediction by showing the combined positive and negative contributions of each feature. For this particular prediction, the most impactful positive features are and , which push the prediction upwards towards 0.84. Conversely, features such as , , and have a slight negative effect, reducing the prediction. The plot effectively highlights which factors led to the final prediction and in which direction they influenced the outcome. Figure 12(b) summarizes multiple predictions, revealing the cumulative effect of features across different samples. For this group of predictions, and emerge as highly influential features. Other features like and also play significant roles, but appears repeatedly as a strong driver in either direction. The red areas show that for many instances, and tend to increase the predicted probability of heart disease, while blue areas for push it down for others. This global view emphasizes the consistency of these critical features in shaping the model’s predictions, reflecting their significant role in cardiac risk assessment.
4.3 Conformal prediction analysis
@@
The calibration curve in Figure 13(a) illustrates the relationship between the predicted probabilities of KACQ-DCNN and the actual fraction of positive outcomes. An ideal calibration curve would follow the diagonal line, indicating that the predicted probabilities align perfectly with the true probabilities. However, as observed, the model exhibits miscalibration in several regions. At lower-mid predicted probabilities (30%–60%), the model tends to underestimate the occurrence of positive outcomes, with the curve falling below the ideal line. Conversely, in the upper-mid range of predicted probabilities (60%–80%), the model overestimates the likelihood of positive outcomes, as indicated by the curve exceeding the diagonal. This overestimation can result in a greater number of FPs. At lower (0%–30%) and higher probabilities (80%–100%), the model’s predictions are closer to the actual outcomes, although slight overprediction remains. These deviations indicate that the predicted probabilities do not fully represent the true likelihoods, especially in the lower-mid and upper-mid ranges. To address this issue, recalibration methods such as Platt Scaling or Isotonic Regression could be employed to improve the alignment between predicted probabilities and actual outcomes.
@@
Figure 13(c) shows that the Standard conformal predictor (left plot) deviates from the diagonal line, mainly in the middle range of p-values. This suggests that it might not be as confident in its predictions for some ranges of positive class probabilities. Similarly, the Mondrian conformal predictor (right plot) also demonstrates deviations from the ideal line. However, the nature and magnitude of these deviations differ because of the additional conditioning on feature bins (Mondrian approach). The Standard and Mondrian conformal predictors initialized via “crepes” were tested on several important metrics, including error rate, average prediction set size, single-class predictions, and number of empty sets. In addition, we recorded the computational efficiency with respect to fitting and evaluation time. These results provide insights into the calibration, precision, and computational feasibility of the model.
@@
We developed a custom wrapper to integrate a TensorFlow-based sequential model with the MAPIE framework for conformal predictions. Early stopping was applied on the basis of validation loss to prevent overfitting. We generated class probability estimates by leveraging the model’s outputs, which were then used to form class predictions based on a threshold of 0.5. This TensorFlow-wrapped model was then integrated into MAPIE’s conformal prediction framework to estimate prediction sets with a confidence level . Specifically, we use sigmoid outputs from the wrapper model as conformity scores, and MAPIE computes the conformal quantile , which helps define the prediction sets for new data points. These prediction sets were designed to include all classes whose sigmoid outputs exceeded , ensuring that the true label was included in the set with a specified probability. Finally, we evaluated the model via different methods (‘score’ and ‘lac’). We tested multiple confidence levels () to generate prediction intervals, ensuring the inclusion of true labels with high reliability across test datasets. Figure 13(b) shows that the estimated quantile values decreased as increased. A higher results in a lower quantile threshold, including more data points in the prediction set. In contrast, a lower (e.g., ) leads to a higher quantile threshold, allowing fewer but more confident predictions. Consequently, when is exceptionally high, it may lead to a low quantile value, resulting in large prediction regions. Conversely, if is too small, the quantile increases, potentially leaving no null areas of the prediction set.
@@
As shown in Figure 14(a), when class coverage is low, there may be times when the model is unsure about its predictions, especially close to decision boundaries. This can cause prediction sets to be empty. As class coverage increases, these empty regions disappear, but areas of ambiguity emerge where both classes may be included in the prediction sets. This means that the selection of class coverage directly influences the prediction decisions and vice versa. A notable situation arises when predictions are based on a threshold of 0.5, as in our binary classification. In this case, each prediction set contained only a single class, eliminating ambiguous or uncertain regions. Figure 14(c) shows that both methods provide coverage close to the target coverage, regardless of the value. The average size of prediction sets increased, and the number of empty prediction sets decreased sharply as increased. The histograms in Figure 14(b) help visualize how often each method produces single or multiple class predictions. The bars for size = 0 are insignificant, implying that the methods do not struggle with uncertainty. Similarly, a dominance of size = 1 indicates strong model confidence.
4.4 emissions in our research
@@
The development, training, and comprehensive evaluation of our proposed KACQ-DCNN model and the associated benchmark models (as detailed in Sections 3 and 4) were computationally intensive. These experiments were carried out on the Google Cloud Platform, specifically in the “asia-east1” region, with a carbon efficiency rating of 0.56 kg of equivalent per kilowatt-hour (kWh). The research conducted hours of computation using T4 hardware, which has a thermal design power (TDP) of 70W. The estimated total emissions were 15.68 kg of CO2 equivalent, all fully offset by the cloud service provider. To clarify the calculations, the power consumption was determined by multiplying the power rating (70W) by the computation time (400 hours), resulting in a total energy consumption of 28 kWh. This figure was then multiplied by the region’s carbon efficiency of 0.56 kg eq/kWh to yield the total emissions of 15.68 kg eq. If the same model was executed in the Google Cloud Platform’s “europe-west6” region, the carbon emissions would have been lower, at 0.45 kg of CO2 equivalent per kWh.
@@
4.5 Interpretation of Results
Our ablation study highlights the synergistic effects of the hybrid approach in the KACQ-DCNN model. While the ROC-AUC column shows significant variance in the ablation test, the KAN version of KACQ-DCNN 1-Layered outperforms all metrics when every component works in harmony. Model-agnostic XAI, such as LIME and SHAP, helped us precisely identify key features affecting predictions, improving the model’s local and global interpretability and trustworthiness by validating feature contributions and ensuring consistent decisions across different instances. These interpretability measures provide confidence in the model’s decision-making process, which is crucial for medical applications where understanding the rationale behind predictions is as important as the predictions themselves.
4.6 Implications
@@
The KACQ-DCNN model demonstrates the potential of classical-quantum hybrid architectures in medical diagnostics, representing a significant advancement in computational approaches to heart disease detection. Uncertainty quantification analysis provided vital insights into the predictive performance and reliability of the model. The calibration curve indicated notable miscalibration in specific probability ranges, suggesting that recalibration methods could improve the model’s probability estimates. The Mondrian approach enhanced reliability with a lower error rate and smaller prediction set sizes, leading to more accurate predictions when used for conformal prediction analysis. Additionally, the Mondrian method demonstrated higher confidence by producing single-class predictions in more cases, further validating the robustness of our approach and its potential for clinical implementation.
4.7 Limitations
Despite promising results, our approach faces several limitations that warrant consideration. The dependency on quantum hardware availability, which is still in the early development stages, limits the practical implementation and scalability of the KACQ-DCNN model. Quantum decoherence issues present technical challenges, as qubits can lose entanglement from heat and light, causing data loss that could affect diagnostic accuracy. The potential for incorrect qubit rotations leading to inaccurate results raises concerns about possible misdiagnoses in clinical settings. Current CQML models are limited by quantum errors, affecting reliability in sensitive applications where precision is paramount. The complexity of model training, requiring careful tuning of classical and quantum components, presents implementation challenges for widespread adoption. Furthermore, the model’s sensitivity to data quality and representativeness, particularly with short and tabular datasets, could impact performance in diverse clinical environments. Finally, the dual-channel design choice may diminish the architecture’s effectiveness in real-time scenarios, potentially limiting its utility in emergency medical situations that require immediate diagnostic feedback.
4.8 Future Work
To further improve model accuracy, scalability, and interpretability, future research directions can explore several avenues. One promising trajectory lies in leveraging quantum advantage through fault-tolerant quantum computing, which could unlock deeper optimization of dressed quantum circuits beyond current noise thresholds. Future models may integrate variational quantum kernels, allowing more expressive decision boundaries suitable for non-Euclidean data structures such as graphs or time series. Integration with symbolic AI could enable KANs to learn interpretable symbolic rules directly from data, bridging the gap between connectionist and rule-based systems. Additionally, coupling KANs with causal inference frameworks could aid in deriving explainable causal relationships from medical datasets. In terms of scalability and generalization, federated learning approaches incorporating privacy-preserving quantum federated KACQ architectures may enable decentralized yet secure training across hospital networks. Similarly, using self-supervised or contrastive learning pretraining in hybrid KAN-QML models could minimize the dependency on labeled data while preserving interpretability. Moreover, incorporating neuro-symbolic reasoning and human-in-the-loop feedback mechanisms into KACQ-DCNN could facilitate real-time clinical collaboration and decision justification, ensuring both trustworthiness and adaptability in high-stakes domains like healthcare.
5 Conclusion
In this study, we proposed the KACQ-DCNN model, a novel classical-quantum hybrid architecture that replaces MLP with KAN layers for heart disease diagnosis. Our comprehensive benchmarking revealed that the KAN-enhanced KACQ-DCNN 1-Layered model outperformed 34 benchmark models, including 16 CML models, 12 QNN models, and 6 hybrid models. The model achieved an accuracy of 92.03% 0.19% with strong performance across multiple metrics, including maP (92.00% 0.19%), maR (92.00% 0.10%), maF1 (92.00% 0.07%), ROC-AUC (94.77% 0.43%), MCC (83.85% 0.81%), and Cohen’s kappa (83.84% 0.43%). Statistical significance was established via the Bonferroni correction (), with all p-values falling well below this threshold, confirming KACQ-DCNN’s superiority.
Future research should aim to incorporate more diverse and representative biomedical datasets to improve the generalizability of the KACQ-DCNN model across varying patient populations and clinical environments. Further advancements in uncertainty quantification techniques could enhance the reliability and clinical trustworthiness of predictive outcomes. Refining the model architecture with a focus on interpretability may facilitate its adoption by healthcare professionals, enabling more transparent decision support. As quantum computing technologies advance, subsequent studies may leverage next-generation quantum systems to achieve increased computational efficiency and scalability. Finally, expanding the diagnostic scope of KACQ-DCNN to include related conditions such as diabetes and sleep apnea represents a critical step toward developing a unified and comprehensive tool for multi-disease screening and diagnosis in clinical practice.
@@ 5 Conclusion
The datasets used in this study are publicly available on Kaggle at the following link:
https://www.kaggle.com/datasets/fedesoriano/heart-failure-prediction.